### v2

# A Theoretical Analysis of Detecting Large Model-Generated Time Series

Abstract
Motivated by the increasing risks of data misuse and fabrication, we investigate the problem of identifying synthetic time series generated by Time-Series Large Models (TSLMs) in this work. While there is extensive research on detecting model generated text, we find that these existing methods are not applicable to time series data due to the fundamental modality difference, as time series usually have lower information density and smoother probability distributions than text data, which limit the discriminative power of token-based detectors. To address this issue, we examine the subtle distributional differences between real and model-generated time series and propose contraction hypothesis, which states that model-generated time series, unlike real ones, exhibit progressively decreasing uncertainty under recursive forecasting. We formally prove this hypothesis under theoretical assumptions on model behavior and time series structure. Model-generated time series exhibit progressively concentrated distributions under recursive forecasting, leading to uncertainty contraction. We provide empirical validation of the hypothesis across diverse datasets. Building on this insight, we introduce the Uncertainty Contraction Estimator (UCE), a white-box detector that aggregates uncertainty metrics over successive prefixes to identify TSLM‑generated time series. Extensive experiments on $32$ datasets show that UCE consistently outperforms state-of-the-art baselines, offering a reliable and generalizable solution for detecting model-generated time series.
1 Introduction
Recent advances in time series forecasting have given rise to Time Series Large Models (TSLMs), which are pre-trained on massive multi-domain time series datasets with billions of parameters (Chronos; timer; TimeMoE). The vast training data and enormous parameter scale enable TSLMs to achieve remarkable long-term zero-shot forecasting on entirely new datasets or domains without any labeled examples or task-specific fine-tuning, as evidenced by recent scaling law analyses (scalinglaw). Leveraging these sophisticated forecasting capabilities, TSLMs have demonstrated strong performance in domains such as finance, Internet of Things (IoT), and climate science.
The powerful capability of TSLMs raises significant concerns about potential data fabrication or misuse. Unlike classical approaches (e.g., ARIMA, LSTM), which degrade in performance outside their training domains, TSLMs can generate coherent long sequences even for unfamiliar domains. If maliciously exploited, this capability could enable the systematic fabrication of time series and pose severe threats in scenarios where data authenticity is critical.
  • Finance: TSLMs can synthesize long transaction histories that mirror real trading patterns, facilitating fraudulent activities such as inflated valuations or hidden manipulations, such as those seen in the 2012 LIBOR scandal (gupta2024defending; rose2013barclays).
  • Scientific Research: Highly realistic counterfeit measurement time series (e.g., signal traces or biological data) can distort experimental outcomes, similar to Schön and Wakefield data forgeries (Bellfabric; godlee2011wakefield).
  • Environmental Governance: Attackers can generate counterfeit metrics (e.g., air quality indices or emissions) that reproduce genuine diurnal and seasonal cycles to conceal pollution spikes or overstate improvements, misleading policy makers and obscuring real hazards (Wang2021).
Figure 1: Illustration of the variation in uncertainty for real and model-generated time series.
To address these threats, we introduce a theoretical framework for white-box detection of TSLM-generated time series. Since there are no dedicated detection methods for time series data, we adapt text-based detection methods to time series data. Textual detectors exploit the observation that LLMs exhibit different token-level probability patterns on human-written and model-generated text, and therefore typically build zero-shot classifiers from each token’s probability or its rank within the model’s vocabulary. However, these methods face significant challenges when applied to time series because of fundamental modality differences.
Textual data exhibit structural semantics and carry rich token-level information (textInformation), which creates greater semantic distances among different tokens. In any given context, only a small subset of tokens is semantically plausible and with higher probability (e.g., probabilities might concentrate on “apple”,“orange” given the prefix “I eat an”). This produces sharper, low-entropy probability distributions over the vocabulary, with a narrower range of token selection.
In contrast, time series carry less information at each point (64words) and inherently contain greater intrinsic uncertainty, yielding smoother probability distributions. Notably, despite the greater entropy of these smooth distributions, they reflect the information at each time point rather than specific values. Since the adjacent values in time series are highly similar (e.g. temperature $25.1^{\circ}\mathrm{C}$ and $25.2^{\circ}\mathrm{C}$), resulting in large mutual information, time series values convey less information. Consequently, text-based detectors that rely on token probabilities perform poorly on time series given the relatively lower probability gaps between values.
Since point-wise probabilities are insufficiently discriminative in time series, we instead analyze the full distributions from the model. We show that a TSLM’s internal distributions, conditioned on the full history, accurately capture the true series distributions and their inherent uncertainty. As TSLMs are trained to minimize prediction errors, the internal distributions are concentrated via model’s sampling strategies for forecasting. In recursive forecasting, the uncertainty is cumulatively decreasing, leading to progressively concentrated internal distributions, as illustrated in Fig. 1.
We therefore propose the contraction hypothesis, namely, TSLM-generated time series exhibit progressively decreasing uncertainty, whereas real time series do not. To validate this hypothesis, we provide a theoretical analysis under idealized assumptions on model behavior and time series structure (see Section 4.2) with empirical evidence through long-horizon forecasting experiments (see Fig. 2). Grounded in this detailed analysis of properties unique to time series, we introduce the Uncertainty Contraction Estimator (UCE), a white-box model generation detection method for time series data. UCE captures uncertainty dynamics from internal prediction distributions and identifies sequences with lower uncertainty levels as model-generated.
(a) Entropy
(b) Max-Probability
(c) Variance
Figure 2: The empirical results show the trajectories of uncertainty metrics, including entropy (2a), max-probability (2b) and variance (2c) of both real and model-generated time series data, illustrating reduction in uncertainty for generated data.
Our main contributions are summarized as follows.
  • To the best of our knowledge, we present the first framework for white-box detection of TSLM-generated time series. We analyze the detailed properties unique to time series in contrast to textual data and address the challenge of low information density in time series.
  • We propose the contraction hypothesis, which states that model-generated time series exhibit progressively decreasing uncertainty during recursive forecasting, whereas real series do not. We provide a theoretical analysis under idealized assumptions of time series and model and empirically validate it.
  • Based on this hypothesis, we develop the Uncertainty Contraction Estimator (UCE), which captures uncertainty dynamics over successive prefixes using TSLMs to distinguish model-generated from real time series.
2 Related Work
2.1 Time Series Large Models
Time series large models have emerged to significantly advance zero-shot time series forecasting. Earlier methods such as PromptCast (PromptCast) and LLMTime (Time_Series_Forecasters) directly leverage LLMs for time series. They convert time series data into text-based prompts and use pretrained LLMs with little to no task-specific adaptation. Such methods require dataset-specific templates and rely heavily on model scale.
PatchTST (64words) introduces patchifying time series into embeddings, a technique later adopted in models such as GPT4TS (OneFitsAll) and Time-LLM (TimeLLM). These methods use patch embeddings as inputs to LLMs for prediction, but require extensive fine-tuning. Moreover, their performance and inference efficiency have been questioned (AreLLMUseful).
Recent work concentrates on models trained on vast and diverse time series datasets. MOMENT (moment) employs uniform random masking to generate patch embeddings for self-supervised pretraining. Chronos (Chronos) discretizes real-valued time series through scaling and quantization for forecasting and introduces data augmentation to improve generalization. Timer (timer) adopts a decoder-only transformer architecture, enabling multiple temporal tasks such as forecasting, anomaly detection, and imputation. Time-MoE (TimeMoE) leverages a Mixture-of-Experts framework to improve forecasting performance and enhance cross-domain generalization while maintaining computational efficiency. These methods collectively represent the emerging trend of general-purpose time series models capable of handling diverse tasks with minimal or no fine-tuning.
2.2 Model Generation Detection
The growing generative capabilities of LLMs have raised concerns about distinguishing model-generated content, particularly in the textual domain. Early work on model generation detection conducts supervised classification using bag of words (Release_Strategies) or neural representations (automatic; authorship), but often overfit and underperform on out-of-distribution data (deepfaketext). To overcome this limitation, zero-shot detectors use LLM output statistics, such as perplexity (relative_entropy) and log rank (gltr; unifying).
Recent work analyzes token probabilities and uses the discrepancies between human-written and model-generated text to develop classifiers. DetectGPT (detectgpt), Fast-DetectGPT (fast_detectgpt), and NPR (detectllm) compare probability differences between perturbed texts, while DNA-GPT (dnagpt) leverages regeneration to compute divergence. FourierGPT (FourierGPT) performs a spectral analysis on token probability sequences to extract linguistic features that distinguish human-authored from model-generated text. Binocular (Binocular) leverages both an observer and a performer to compute cross-perplexity and isolate intrinsic prefix-induced uncertainty to improve detection accuracy. Black-box methods such as intrinsic dimension (Intrinsic_dimension) perform topological data analysis over token embeddings and find that human-written texts are more complicated in topology.
Despite their success in text, these methods rely on the dense and discrete structure of language, making them unsuitable for continuous and information-sparse time series data. In this work, we extend zero-shot model generation detection to time series by introducing uncertainty-based metrics derived from model’s internal probability distributions.
3 Problem Formulation and Preliminaries
In this section, we formally define the model-generated time series detection problem. Then we provide some preliminaries about our framework.
3.1 Problem Formulation
Let $\mathbf{X}_{t}=(X_{1},\ldots,X_{t})$ denote a univariate time series with unknown history. We perform a zero-shot classification of $\mathbf{X}_{t}$ as real or model-generated without labeled examples.
We assume a white-box setting with access to the model’s internal probability distribution $p_{\theta}(\cdot|\mathbf{X}_{t})$ rather than the model architecture or parameters $\theta$. Our detection employs point-wise probabilistic TSLMs over an equidistant vocabulary $\mathcal{V}=\{v_{i}|v_{i+1}-v_{i}=\Delta,-R\leq v_{i}\leq R\}$ for new token sampling. Therefore, sampling token $v_{i}$ yields the numerical forecast $\hat{X}_{t+1}=v_{i}$. In the limit $\Delta\to 0$ and $R\to\infty$, $\mathcal{V}$ converges to $\mathbb{R}$ and induces a probability density $f_{\theta}$. Such a formulation allows for detection evaluation across diverse model architectures.
Some frequently used symbols throughout the paper are given in Table 1. In the following sections, we refer to distributions by their probability densities.
Symbol Description
${X}_{t},\hat{X}_{t}$ Real / forecast time series process
$T_{t}$ Deterministic trend component at time $t$
$n_{t},\hat{n}_{t}$ True / forecast (Gaussian) noise at time $t$
$\mathbf{X}_{-H:t},\hat{\mathbf{X}}_{-H:t}$ History $X_{-H},\ldots,X_{t}$ (or $\hat{X}_{-H},\ldots,\hat{X}_{t}$)
$f_{t}(X_{t}),\hat{f}_{\theta}(\hat{X}_{t})$ Probability density of $X_{t}$ / $\hat{X}_{t}$ at time $t$
$\sigma_{t}^{2},\hat{\sigma}_{t}^{2}$ Variance of true noise $n_{t}$ / forecast noise $\hat{n}_{t}$
$f_{\theta}(z_{t}|\mathbf{X}_{-H:t-1})$ Model’s internal probability density with history $\mathbf{X}_{-H:t-1}$, abbreviated as $f_{\theta}(z_{t})$
$\tilde{\sigma}_{t}^{2}$ Variance of internal probability distribution at time $t$
$\gamma_{t}$ Scaling factor of sampling strategy to $f_{\theta}(z_{t})$ at time $t$
Table 1: Some frequently used symbols
3.2 Preliminaries
We first introduce some assumptions and premises about time series and idealized TSLMs. Detailed definitions and assumptions are provided in the Appendix.
We formally model time series as realizations of stochastic processes, where each observation sequence is a sample path of this process. We decompose the real time series process at time point $t$ into two components (TimeSeriesAnalysis):
$X_{t}=T_{t}+n_{t},\quad n_{t}\sim\mathcal{N}\left(0,\sigma_{t}^{2}\right),$
where $T_{t}$ denotes the trend sequence, representing predictable components, and $n_{t}$ is a Gaussian process with zero mean and variance such that $\sigma_{t}^{2}=\sum_{i=1}^{l}\alpha_{i}\sigma_{t-i}^{2}$, $\sum_{i=1}^{l}\alpha_{i}=1$. The noise process represents the unpredictable or uncertainty components.
We introduce Ideal Model, which, for any real-valued time series in $\mathbb{R}$, predicts probability distributions to minimize the expectation of cross-entropy loss. Conceptually, it generalizes practical TSLMs by letting both training data and model capacity grow without bound. This abstraction allows us to derive fundamental detection principles regardless of specific architectures or training constraints.
We formalize model evaluation function as Eq. 1:
$\mathrm{Eva}=\left[\sum_{i=1}^{\tau}g\left(\left|X_{i}-\hat{X}_{i}\right|^{p}\right)\right]^{1/p},1\leq p\leq\infty,$ (1)
where $g$ is any non-negative strictly increasing mapping. Specifically, if $g$ is linear, $\mathrm{Eva}$ reduces to the standard $\ell_{p}$ norm of point-wise errors (e.g., RMSE when $p=2$). Since $X_{i}$ and $\hat{X}_{i}$ are variables, $\mathrm{Eva}$ is also a variable. $\mathrm{Eva}$ is a unified form of various model prediction evaluators (e.g., MSE, MAE) which in turn guides the model generation process.
4 Methodology
In this section, we introduce the Uncertainty Contraction Estimator (UCE), a novel time series model generation detection method. We first establish the theoretical foundation by proposing the contraction hypothesis in Section 4.1, followed by a formal analysis in Section 4.2. Finally, the implementation details of UCE are presented in Section 4.3.
4.1 Contraction Hypothesis
UCE hinges on two observations: (1) the model’s internal distributions faithfully reproduce true series distributions at each time; (2) through sampling strategies, the model systematically modifies these distributions, creating a self-reinforcing process of exponentially decreasing uncertainty. We formalize these insights as the contraction hypothesis.
Contraction hypothesis: TSLM-generated time series exhibit progressively concentrating internal distributions with decreasing uncertainty, whereas real series do not.
The term “contraction” refers to the contraction mapping nature of the uncertainty reduction. We systematically substantiate the hypothesis through a tripartite theoretical framework. Formal proofs of all the following propositions are provided in the Appendix.
4.2 Theoretical Analysis
Distributional Consistency
We begin with the property of ideal models. Given any history $\mathbf{X}_{-H:t-1}$, TSLM generates an internal probability distribution $f_{\theta}$. From the perspective of this distribution, we revisit the ideal models and conclude the distribution-perfect prediction property. {lemma} For $\sigma_{t}^{2}\geq 0$, we have $f_{\theta}\equiv f_{t}\;a.e.$ Fig. 3 illustrates the insight of Section 4.2. Since the probability density of $X_{t}=T_{t}+d$ is $f_{t}(T_{t}+d)$, the loss expectation is therefore the cross entropy $\int f_{t}(v)\log f_{\theta}(v)\mathrm{d}v$ between $f_{\theta}$ and the true series distribution $f_{t}$. As illustrated in Fig. 3, it is minimized when $f_{\theta}\equiv f_{t}$ almost everywhere by Gibbs’ inequality. Specifically, if the true sequence is entirely deterministic (i.e., zero uncertainty), ideal models make value-perfect predictions, leading to Section 4.2. {corollary}[] If $\sigma_{t}^{2}=0$, we have $f_{\theta}(z=v)=\delta(v-T_{t})$, where $\delta$ is the Dirac-$\delta$ function.
In summary, the ideal model’s internal distributions coincide with true series distributions and therefore preserve the inherent uncertainty in the series.
Figure 3: Comparison of the true distribution (blue) and model’s internal distributions (orange). For an ideal model, its internal distribution coincides with the true distribution.
Sampling-Induced Variance Scaling
Using the property of ideal models to reproduce true series distributions, we analyze the prediction of $\mathbf{\hat{X}}_{1:\tau}$ with history $\mathbf{X}_{-H:0}$, and focus on the variation in uncertainty. In this section, we explain that the sampling probability distribution $\hat{f}_{t}$ of each $X_{t}$ is a modified version of $f_{\theta}$, which tends to reduce its uncertainty.
TSLMs modify internal distributions $f_{\theta}$ through sampling strategies to modulate output diversity (GPT2). We summarize common sampling strategies as follows: (1) scaling (e.g., temperature sampling); and (2) symmetric truncation (with normalization, e.g., top-$k$). These modifications preserve the means of $f_{\theta}$ but alter their variances, as shown in Fig. 4, generating sampling probability distributions $\hat{f}_{\theta}(\hat{X}_{t}=v)$ to sample $\hat{X}_{t}$ with variance $\hat{\sigma}^{2}_{t}$.
Let the uncertainty (variance) of $f_{\theta}$ be $\tilde{\sigma}^{2}_{t}$ at time $t$, we denote the modified uncertainty by sampling strategies as $\hat{\sigma}^{2}_{t}=\gamma_{t}\cdot\tilde{\sigma}^{2}_{t}$. Intuitively, $\gamma_{t}$ controls uncertainty expansion ($\gamma_{t}>1$), preservation ($\gamma_{t}=1$), or contraction ($\gamma_{t}<1$). Since modifications of sampling strategies preserve the expectation, we recursively obtain $\mathbb{E}(\hat{X}_{t})=T_{t}$ for all $1\leq t\leq\tau$. We therefore treat the generation of $\hat{\mathbf{X}}_{1:\tau}$ as $\tau$ independent, one-step predictions given their respective histories. We evaluate the differences between $\mathbf{X}_{1:\tau}$ and $\hat{\mathbf{X}}_{1:\tau}$ using the model evaluation function, leading to Section 4.2. {lemma}[] The expectation of the model evaluation function $\mathbb{E}(\mathrm{Eva})$ increases monotonically with respect to all $\hat{\sigma}_{t}$. Section 4.2 indicates that a smaller $\hat{\sigma}_{t}^{2}$ yields a lower $\mathrm{Eva}$ value, suggesting a higher quality of generation. Given $\hat{\sigma}_{t}^{2}=\gamma_{t}\cdot\tilde{\sigma}_{t}^{2}$, when $\tilde{\sigma}_{t}^{2}$ is fixed, a smaller $\gamma_{t}$ yields lower $\mathbb{E}(\mathrm{Eva})$, which validates our intuition.
Figure 4: Comparison of the internal distribution (blue, from the model logits) and modified sampling distribution (orange, from 10,000-step Monte Carlo) for a single prediction step. The model’s internal probability distribution becomes sharper under our sampling strategy.
In particular, we derive a corollary from Section 4.2 that if metrics such as MSE are used as training loss, the contraction of uncertainty is immediate. {corollary} If the loss function is defined in the same form as the model evaluation function $\mathrm{Eva}$ in Eq. 1 (e.g., MSE),
we have $f_{\theta}(z=v)=\delta(v-T_{t})$.
Based on the analysis, we intuitively expect that TSLMs employ $\gamma_{t}<1$ in practice. In the following, we further analyze this intuition of $\gamma_{t}$ in terms of performance.
Recursive Variance Reduction
With the intuition of $\gamma_{t}$ above, we examine how the sampling distribution (i.e., the internal distribution modified by $\gamma_{t}$) affects the subsequent generation. In this section, we demonstrate the recursive evolution of the variance and analyze its behavior under different values of the scaling factor $\gamma_{t}$.
According to Section 4.2, the ideal model reproduces any distribution with variance satisfying $\sigma_{t}^{2}=\sum_{i=1}^{l}\alpha_{i}\sigma_{t-i}^{2}$. When forecasts $\hat{X}_{1},\hat{X}_{2},\ldots$ serve as histories, their variances $\hat{\sigma}_{1}^{2},\hat{\sigma}_{2}^{2},\ldots$ also follow the same recurrence relation.
$\tilde{\sigma}_{t}^{2}=\sum_{i=1}^{l}\alpha_{i}\hat{\sigma}_{t-i}^{2}=\sum_{i=1}^{l}\alpha_{i}\gamma_{t-i}\tilde{\sigma}_{t-i}^{2},\;\sum_{i=1}^{l}\alpha_{i}=1.$ (2)
Typically, the uncertainty scaling behavior is consistent and $\gamma_{t}<1$ or $\gamma_{t}>1$ for all $t$. When $\gamma_{t}<1$, the internal uncertainty, according to Eq. 2, is strictly lower than the weighted sum of historical uncertainties, leading to an exponential decay in uncertainty towards 0. By Section 4.2, the forecast series $\hat{X}_{t}$ converges to the trend $T_{t}$. Conversely, when $\gamma_{t}>1$, the uncertainty grows exponentially towards infinity, introducing excessive noise into the forecast that undermines trend capture and violates forecasting objectives. When $\gamma_{t}\equiv 1$, the uncertainty in forecast series is relatively stable for ideal models, which is unattainable for practical models due to noise accumulation.
This theoretical analysis also helps explain why many practical TSLMs exhibit similar behaviors of uncertainty contraction. For example, Chronos (Chronos) adopts contractive sampling strategies by using top-$k$ with median sampling, and therefore $\gamma_{t}<1$ holds consistently. Timer (timer) and Time-MoE (TimeMoE) utilize point-wise errors (e.g., MSE) as training loss, sharing the form as $\mathrm{Eva}$ defined in Eq. 1. According to Fig. 4, the forecast uncertainty converges rapidly, indicating their equivalence to $\gamma_{t}<1$ in terms of performance. We formalize the aforementioned analysis as contraction hypothesis, where variance can be substituted for other metrics. {theorem} For an Ideal Model’s forecast time series $\mathbf{\hat{X}}_{t}$ with history $\mathbf{X}_{-H:0}$, let the internal probability distribution at $t$ be $f_{\theta}(z_{t})$, and define its uncertainty measure by variance $\tilde{\sigma}_{t}^{2}$. Therefore we have that $\tilde{\sigma}_{t}^{2}$ is monotonically decreasing with $t$ and $\lim_{t\to\infty}\tilde{\sigma}_{t}^{2}\approx 0$.
We provide empirical verification of Section 4.2 in Figure 2, which illustrates this uncertainty contrast over 1024 tokens using three metrics, namely entropy, max-probability, and variance, which are used in our detection method. In model-generated series, max-probability steadily approaches $1$, while both entropy and variance decay towards 0. In contrast, real sequences remain comparatively stable in uncertainty. This observed discrepancy in series uncertainty, combined with the ability of TSLMs to capture it, forms the basis of our detection methodology.
Figure 5: Average AUROC and TPR (at 1% FPR) for model generation detection on In-Distribution (12 datasets) and Zero-Shot (20 datasets) scenarios.
4.3 Uncertainty Contraction Estimator
The Uncertainty Contraction Estimator (UCE) operationalizes the Contraction Hypothesis by quantifying the uncertainty contraction in model-generated time series. Given a candidate $\mathbf{X}_{t}$, we first sample $N$ time points $t_{1},t_{2},\ldots,t_{N}$ with a fixed interval $\Delta t=t_{i+1}-t_{i}$. For each $t_{i}$, UCE takes $\mathbf{X}{t_{i}}$ as the input to the TSLM and computes its internal probability distribution $\hat{P}_{t_{i}}$ as defined in Eq. 3, yielding a sequence of internal distributions.
$\hat{P}_{t_{i}}=p_{\theta}\left(\cdot|X_{1},\ldots,X_{t_{i}}\right),i=1,\ldots,N.$ (3)
By Section 4.2 the internal probability distribution coincides with the underlying Gaussian noise distribution, which is unimodal and relatively probability-concentrated. Therefore, UCE focuses on a neighborhood $\mathcal{U}$ around the mean. Within $\mathcal{U}$, UCE computes three uncertainty measures to capture different aspects of the distribution concentration.
  1. 1.
    Entropy $E=-\sum_{x\in\mathcal{U}}\hat{P}\left(x\right)\log\hat{P}\left(x\right)$, which captures the spread of probability mass.
  2. 2.
    Max-Probability $P_{\max}=\max_{x\in\mathcal{U}}\hat{P}(x)$, where larger values imply lower uncertainty.
  3. 3.
    Variance $\mathrm{Var}=\sum_{x\in\mathcal{U}}\left(x-\mu\right)^{2}\hat{P}(x)$ with $\mu$ the local mean to quantify concentration.
For a selected metric $s\in\{E,P_{\max},\mathrm{Var}\}$, UCE calculates the metric sequence $s_{t_{1}},s_{t_{2}},\ldots,s_{t_{N}}$, and the overall UCE score is their mean, formulated as $\mathrm{UCE}=\frac{1}{N}\sum_{i=1}^{N}s_{t_{i}}$.
5 Experiments
5.1 Experimental Settings
Model and Datasets
We evaluate all detection methods using Chronos-T5 (large), a point-forecasting TSLM that discretizes continuous values over a fixed token vocabulary and produces probability distributions. We conduct experiments on 32 datasets across diverse domains (energy, finance, transportation), categorized as (1) in-distribution datasets, including 12 used during Chronos training; and (2) zero-shot datasets, consisting of 20 previously unseen datasets. The zero-shot setting evaluates the generalization of detection to unseen datasets to eliminate the effects of data leakage, and validates the “black-box” detection capability (see Section 5.3). For each dataset, we follow the default setup in Chronos to generate forecasts of horizon $H=64$, and use the corresponding $H$ observations as ground truth. We treat forecast series as positive (model-generated) samples and real series as negative samples.
Baselines
We compare UCE with 10 text-based baselines adapted for time series: (1) DNA‑GPT WScore (dnagpt), which measures the average log‑likelihood gap between regenerated and original text given a fixed prefix; (2) DetectGPT (detectgpt) and its efficient variant Fast‑DetectGPT (fast_detectgpt), which detect generation by measuring probability changes under model‑guided perturbations; (3) DetectLLM (detectllm), using Log‑Likelihood Log‑Rank Ratio (LRR) and Normalized Perturbed log‑Rank (NPR) metrics; (4) Intrinsic Dimension (Intrinsic_dimension), which classifies sequences by estimating the topological “intrinsic dimension” of token embeddings; (5) FourierGPT (FourierGPT), performing spectral analysis on token probability sequences; (6) Binocular (Binocular), leveraging two models to reduce the intrinsic perplexity of prefixes; and (7) Traditional metrics (gltr; automatic; Release_Strategies), including log‑likelihood $\log p(x)$, rank, and log‑rank. The details of the baselines are documented in the Appendix.
Evaluation Metrics
We use the Area Under the ROC Curve (AUROC) to evaluate the detection performance of UCE and the baseline methods. Since AUROC masks performance at low false positive rates (FPR), which are critical in generation detection tasks (1stprinc; paraphrasing; dnagpt), we also report the true positive rate (TPR) at a fixed FPR of 1% (curiouscaseAUROC).
5.2 Overall Result
We evaluate detection performance under in-distribution and zero-shot settings. UCE consistently achieves state-of-the-art AUROC and TPR across all scenarios, with key results shown in Fig. 5.
In-Distribution
UCE-Entropy achieves an AUROC of 0.855, outperforming the strongest baseline, DetectLLM-LLR (0.815), by 0.040, and exceeding the baseline average (0.670) by 0.183. Its TPR reaches 0.447, surpassing DetectLLM-LLR (0.324) by 0.123 and the baseline average (0.118) by 0.329.
Zero-Shot
In the zero-shot setting, UCE-Entropy achieves an AUROC of 0.731, outperforming DetectLLM-LLR (0.705) by 0.026 and the baseline average (0.632) by 0.099. It also attains a TPR of 0.286, exceeding DetectLLM-LLR (0.233) by 0.053 and the average baseline (0.151) by 0.135.
The other two UCE metrics (Max-Probability and Variance) also show strong performance, with UCE–MaxProb approaching UCE–Entropy in TPR despite a modest AUROC gap; UCE–Variance is slightly weaker in TPR. Overall, UCE-Entropy remains the top performer, UCE-MaxProb is a close second, and UCE-Variance may offer greater robustness under alternative conditions. The full experimental results, with additional experiments, are provided in the Appendix. The experimental results suggest potential future work on hybrid or adaptive variant selection.
5.3 Cross-Model Detection Performance of UCE
In this section, we evaluate the generalization of UCE on time series generated by two alternative TSLMs, Timer and Time-MoE. We conduct experiments on 9 datasets covering multiple lengths and report both AUROC and TPR.
UCE achieves strong overall performance on both models, with the Entropy variant exhibiting particularly notable results, consistently achieving high AUROC and TPR across both models, as shown in Table 2 and Table 3. The Max-Probability and Variance variants yield weaker performance, especially for Timer-generated sequences. Notably, UCE performs particularly well in Time-MoE, possibly due to its Mixture-of-Experts architecture with better long-forecasting performance. In Section 4.2 we show that although Timer and Time-MoE are not probabilistic forecasting models, they also exhibit uncertainty contraction, and uncertainty metrics demonstrate discriminative power between real and model-generated series.
Furthermore, the performance of UCE on both zero-shot datasets and cross-model settings also implies its potential for the “black-box” detection analogous to DetectGPT. Specifically, given a time series that originates from an unknown model and unknown source (or the real world), UCE can perform detection by leveraging a locally deployed probabilistic model to compute uncertainty-based signals.
Time Series
Length $H$
UCE-Entropy UCE-Max Prob UCE-Variance
AUROC
TPR
AUROC
TPR
AUROC
TPR
$H=96$ 0.833 0.301 0.556 0.041 0.635 0.108
$H=192$ 0.771 0.280 0.508 0.050 0.576 0.102
$H=336$ 0.765 0.305 0.484 0.032 0.564 0.093
$H=768$ 0.788 0.366 0.542 0.096 0.602 0.179
Table 2: AUROC and TPR (at 1% FPR) for Timer generation detection on 9 datasets.
Time Series
Length $H$
UCE-Entropy UCE-Max Prob UCE-Variance
AUROC
TPR
AUROC
TPR
AUROC
TPR
$H=96$ 0.829 0.320 0.717 0.292 0.745 0.316
$H=192$ 0.890 0.392 0.773 0.420 0.806 0.422
$H=336$ 0.957 0.611 0.810 0.475 0.856 0.511
$H=720$ 0.950 0.561 0.845 0.540 0.863 0.566
Table 3: AUROC and TPR (at 1% FPR) for Time-MoE generation detection on 9 datasets.
6 Discussion
Our methodology is based on the contraction hypothesis, which depends on certain idealized assumptions (see Section 4.1). Despite the empirical support for the hypothesis from the experimental results, we revisit the foundational assumptions to strengthen our theoretical analysis.
6.1 Idealized Model Assumption
To establish an architecture-agnostic detection methodology, we postulate a theoretically optimal Ideal Model. This abstraction avoids model-specific details to identify general behavior, but is mathematically unrealizable in practice. Existing TSLMs are approximations of this ideal, which undermine both the guarantee of perfect prediction (see Section 4.2) and the model’s ability to faithfully recover the true series distribution (see Section 4.2). Despite the limitation, practical models also exhibit similar uncertainty contraction behavior in forecasting.
In particular, TSLMs are trained on finite datasets with finite parameters. Since the noise samples are limited and the model is trained to minimize expected loss over these few realizations, it cannot fully capture the underlying noise distribution or the true trend. Therefore, both the estimated trend $\hat{T}_{t}$ and noise $\hat{n}_{t}$ inevitably deviate from the truth, and the trend deviation $\Delta T_{t}=|T_{t}-\hat{T}_{t}|$ accumulates as forecasting progresses. Given this non-zero deviation, Section 4.2 does not strictly hold, and reducing forecast variance $\hat{\sigma}_{t}^{2}$ does not always improve evaluation performance. Nevertheless, a moderate reduction in uncertainty constrains the expected evaluation error within the relatively small bound $\Delta T_{t}$ by concentrating probability between $T_{t}$ and $\hat{T}_{t}$. As the deviation grows, greater uncertainty does not yield sufficiently better performance by correcting the error and may result in unstable forecasts. This indicates that practical TSLMs also inherently exhibit a forecasting uncertainty contraction (i.e., $\gamma_{t}<1$), especially under recursive noise accumulation, as elaborated in Section 4.2. Hence, the contraction hypothesis still holds for practical models, allowing UCE to track predictive uncertainty across recursive steps.
6.2 Gaussian Noise Assumption
In this work, we assume Gaussian noise for analytical simplification: the sum of independent Gaussian random variables remains Gaussian, and the variance fully characterizes the distribution. Crucially, our proofs do not depend on the exact form of the Gaussian but only require the noise to be unimodal (greatest probability for trend), symmetric (equivalence for both positive and negative biases), and having a finite second moment (existence of variance). By the additivity of the variance for independent variables, our results naturally extend to such noise models. These properties are shared by most noise models commonly assumed in real-world time series analysis (e.g., Laplace), and the experimental results (see Appendix) validate the broad applicability of the contraction hypothesis under various noise types.
6.3 Modality Difference Analysis
In this work, we extend textual detection methods to time series via UCE. Although uncertainty metrics (e.g., entropy) are relatively simple, we investigate the unique properties of time series versus text modalities (see Section 4.2) to prove the optimality of uncertainty in discriminative power.
7 Conclusion
We investigate detecting TSLM-generated time series and hypothesize that they exhibit progressively decreasing uncertainty—unlike real data. Building on this, we propose the Uncertainty Contraction Estimator (UCE), which captures uncertainty to distinguish model-generated from real series and is validated both theoretically and empirically. Future work will further analyze the hypothesis under diverse model architectures and extend UCE to multivariate and batch-forecasting settings.
Acknowledgements
This work was supported in part by National Key Research and Development Plan in China (2023YFC3306100) and National Natural Science Foundation of China (62272372).